Circular No. 6600

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) J. Sarmecanic, University of California at San Diego (UCSD); D. J.

Osip, University of Florida; and M. Fomenkova and B. Jones, UCSD, report: “Imaging of comet C/1995 O1 has been obtained using the UCSD midinfrared camera at the 1.5-m telescope on Mt. Lemmon, Arizona. The inner coma (40 000 km across) was continuously observed for 11.5 hr at 8.7, 10.3, 11.7 and 12.5 μm during Mar. 3.51–4.00 and 5.52–6.00 UT. A bright jet, initially directed to the southwest, appears in all filters; over time, it rotates clockwise and reappears in the southwest 11.3 hr later, consistent with previously reported rotation periods of between 11.2 and 11.65 hr (IAUC 6560, 6583). There are also two arcs on the sunward side (southwest) at 15′′ and 30′′ from the nucleus. Photometric magnitudes in a 10′′ circular aperture on Mar. 5.53, 5.77 and 5.97 are, respectively (uncertainties 0.10–0.15 mag): [8.7 μm] = −4.6, −4.4, −4.4; [10.3 μm] = −6.0, −5.6, −5.6; [11.7 μm] = −5.7, −5.6, −5.8; [12.5 μm] = −5.2, −5.0, −5.1. The strong silicate feature, as observed at 10.3 μm, is a factor of 2.5 ± 0.3 higher than the continuum, which is estimated to be 390 K.”

GRB 920501 M. J. Drinkwater and L. A. Jones, University of New South Wales;

O. M. Blaes, T. Hurt and R. Antonucci, University of California at Santa Barbara; and K. Hurley, University of California at Berkeley, report: “We have measured the optical spectrum of the galaxy in the HRI error circle of the possible x-ray counterpart to GRB 920501 with the Australian National University 2.3 m telescope at Siding Spring Observatory. The galaxy is labelled source 7 in Fig. 1 of Blaes et al. (1997, Ap.J., in press). This figure is available on the World Wide Web at http://vela.physics.ucsb.edu /people/jkeck.html. The HRI error circle depicted there has 6′′ radius and center at α = 8^{h}15^{m}18^{s}.06, δ = −32^{o}44′27′′.0 (equinox 2000.0). The spectrum shows Hβ, [O III] (495.9 and 500.7 nm), Hα and the [S II] doublet in emission at a redshift of 0.315. The Hα line has an equivalent width of ∼ 40 nm and an FWHM of 2300 km/s. This strongly suggests that the galaxy is a type 1 Seyfert and that it is the x-ray source. Given the possible association with the γ-ray burst, we encourage further observations of this Seyfert.”

1997 March 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden